OpenStudy (anonymous):

Can someone please explain to me what a function is? I fan and give medals! :)

OpenStudy (anonymous):

function is a relation.

OpenStudy (anonymous):

I know that the x cant have the same number twice but what about the bottom? (y) can it have the same number?

random231 (random231):

well a function can be defined as a relation between generally two variables. for ex y=x+2 is a function of x where every value of x gives a definite value of y.

OpenStudy (anonymous):

A function is a special relationship between values: Each of its input values gives back exactly one output value. It is often written as "f(x)" where x is the value you give it. Example: f(x) = x/2 ("f of x is x divided by 2") is a function, because for every value of "x" you get another value "x/2", so: * f(2) = 1 * f(16) = 8 * f(-10) = -5

OpenStudy (anonymous):

ok thanks guys!

random231 (random231):

yes it can have the same number.